OpenStudy (anonymous):

Find an equation of the line passing through the point (6,5) and perpendicular to the line whose equation is 2y+3x=6.

OpenStudy (anonymous):

Given equation is 2y+3x=6 = 2y=6-3x =y= 3-3/2x now we have equation of line in a form of y=mx+c slope m=-3/2 prependicular slope=2/3 now equation of line is given by y=mx+b given points are (6,5) plug values we will get 5=2/3(6)+b 5=4+b b=1 now equation of line prependicular to given line is y=2/3x+1

OpenStudy (anonymous):

ty